1. If 8 teachers can mark 450 scripts in 5 days, how many scripts will it take 20 teachers to mark in 8 days?


1
Expert's answer
2021-05-07T12:31:10-0400

One teacher in one day can mark: "\frac{450}{5*8}=11.25" scripts.

20 teachers in 8 days can mark: "11.25*8*20=1800" scripts.